# Diffusion based Semantic Outlier Generation via Nuisance Awareness for Out-of-Distribution Detection

### v1

@@ 1 Introduction
Figure 2: Performance comparison of Near- vs Far-OOD detection with auxiliary outliers. Only OE utilizes real outliers; the rest methods use synthetic. Outliers for the junco bird species (ID) are shown on the left. Around 10% AUROC drop is observed in Near-OOD across all methods.
@@ 3.4 SONA Guidance
In this section, we introduce SONA guidance, a novel approach that enables fine-grained control on each semantic and nuisance region for outlier generation. Our method draws inspiration from recent advances in image editing, which aims to completely replace target semantics. However, as they are not inherently designed for OOD detection, their application has shown limited impact in this context (see Appendix G). We propose a developed approach adequate for outlier generation, allowing precise control over both $M_{\mathtt{S}}$ and $M_{\mathtt{N}}$ to prioritize semantic differences. The SONA guidance term, $\Delta_{\texttt{SONA}}$, is composed of three components as follows, each of which is described in the following paragraphs:
@@
Finally, we obtain the outliers by denoising $\mathbf{z}_{t}$ for every timestep $t=\tilde{T},\dots,1$ with our SONA guidance. To generate the pixel-space outlier image $\mathbf{x}_{\texttt{OOD}}$, we pass the denoised latent representation $\mathbf{z}_{0}$ through the decoder of the diffusion model. Algorithm 2 in the Appendix provides a detailed pseudo-code implementation of our framework.
@@ 3.5 OOD Detection with SONA Outliers
During the test phase, we utilize the energy score (Liu et al. 2020a), which effectively addresses overconfidence issues in OOD detection by assigning lower energy values to ID samples and higher energy values to OOD samples.
@@ 4.1 Experimental Setup
We mainly evaluate our framework on ImageNet-200 (Zhang et al. 2023b) as ID, a subset of 200 categories from ImageNet-1k (Deng et al. 2009). Our evaluation covers both far-OOD and challenging near-OOD scenarios. For far-OOD detection, we employ widely-used datasets such as iNaturalist (Horn et al. 2018), Texture (Cimpoi et al. 2014a), and OpenImage-O (Wang et al. 2022). For near-OOD detection and SSB-hard (Vaze et al. 2022a) and NINCO (Julian Bitterwolf and Hein 2023) for near-OOD detection., which have no class overlap but show close semantic similarity with ImageNet-1K.
@@
Our implementation is based on Stable Diffusion v2-base model  [^1] See https://huggingface.co/stabilityai/stable-diffusion-2-base for more details., with sampling hyperparameters consistent with (Brack et al. 2024). For selecting $\mathbf{c}_{\texttt{OOD}}$, we utilize labels from the remaining 800 classes of ImageNet-1k (disjoint classes from ImageNet-200) for the main table. The number of generated samples is equal to the number of entries in the ID training dataset. For the OOD detector training, we employ ResNet-18 (He et al. 2016) as the network architecture. Additional implementation details can be found in the Appendix D
@@ 4.2 Main Results
Table 1 presents a comprehensive comparison of OOD detection performance on ImageNet-200, highlighting the superiority of our framework. Our proposed method consistently outperforms in both near and far-OOD scenarios, affirming its effectiveness. It surpasses all benchmarks in the near-OOD setting, achieves top results in two out of three far-OOD datasets, and attains the second-best performance in the remaining one. Remarkably, our method achieves a state-of-the-art (SOTA) AUROC of 88.4% on near-OOD detection, showing exceptional performance in this challenging scenario. This score represents a notable improvement, outperforming other synthesis-based methods by 6%. Furthermore, while other baseline methods experience a sharp decline of 10% in near-OOD compared to far-OOD, our approach significantly reduces this performance gap. Another noteworthy point is we even surpass real-world outlier based methods, without requiring any additional datasets. This indicates that our methodology is also cost-effectively feasible for real-world applications. Another superior performance of SONA across a broader range of settings, including full-spectrum (e.g., covariate-shifted ID) and fine-grained benchmarks, can be found in the Appendix  F.
@@
Figure 6: Performance comparison of SONA guidance and global guidance across $\tilde{T}$. (a) AUROC (%) analysis. (b) LPIPS score analysis.
Figure 7: (a) Ablation study on region masking threshold $\lambda$ (b) Ablation study of SONA guidance scale $s$.
@@ 4.3 Analysis
Table 2: Comparison of SONA with different OOD prompt selection.
@@
Table 3: Ablation study on $\Delta_{\texttt{SONA}}$ components.
@@
Table 4: Ablation study on $\mathcal{L}$ components.
@@ 6 Conclusion
Limitation. This paper focuses on extracting and utilizing nuisance information of ID images to generate outliers. However, there exist various types of nuisance information (e.g., drawings styles and textures), and it might require different approaches to extract and utilize them better. We believe our framework can be further specialized to a nuisance type based on additional knowledge (e.g., style representation (Zhang et al. 2023c)) or can be further improved by advanced representation learning (e.g., contrastive learning (Tack et al. 2020)).
@@
Diffusion based Semantic Outlier Generation via Nuisance Awareness
@@ A.2 Classifier-Free Diffusion Guidance
Appendix B Full Algorithm of SONA
@@
Appendix C Dataset Details
C.1 Near-OOD Detection
  • SSB-hard (Vaze et al. 2022b) (Semantic Shift Benchmark - Hard): SSB-hard is a subset derived from the Semantic Shift Benchmark (SSB), designed to evaluate model performance under subtle semantic shifts. It contains images closely similar to ImageNet-1K, making it particularly challenging. The dataset comprises 980 classes from ImageNet-21K, sharing significant visual and semantic characteristics with ImageNet-1K.
C.2 Far-OOD Detection
  • Textures (Cimpoi et al. 2014b): The Textures dataset contains images representing a wide variety of real-world patterns and surface textures, differing from the object-centric images in ImageNet-1K. It tests the model’s ability to recognize OOD samples based on textural differences, crucial for robust OOD detection.
Appendix D Implementation Details of SONA Framework
Implementation details.
Our implementation is based on Stable Diffusion v2-base model  [^2] See https://huggingface.co/stabilityai/stable-diffusion-2-base for more details for more information for generating SONA. For our best results, we set the total number of diffusion steps $T$ to 50, and at that point randomly selected a stop timestep $\tilde{T}$ from the range between 1 and 50. For the OOD text condition, $\textbf{c}_{\texttt{OOD}}$, we use text prompts based on ImageNet-800 words, which do not overlap with ImageNet-200. We set the guidance scale $s$ to 10 and the guidance threshold $\lambda$ to 0.2 for all datasets. We follow the additional hyperparameters for image sampling provided by (Brack et al. 2024). Since we use the ID training dataset directly for SONA generation, the number of SONA is equal to the number of entries in the ID training dataset. While synthesizing our SONA samples, we use 8 A40 gpus. Implementation of SONA is summarized as following algorithm 2:
Baselines.
We compare our framework against existing methods divided into three categories: Post-hoc detection methods that do not require further training, training with real-world outlier methods that require explicit outliers, and outlier synthesis-based methods. For the post-hoc baselines, we compare against OpenMax (Bendale and Boult 2016), MSP (Hendrycks and Gimpel 2017), ODIN (Liang, Li, and Srikant 2018), EBO (Liu et al. 2020b), OpenGAN (Kong and Ramanan 2021), ReAct (Sun, Guo, and Li 2021), KNN (Sun et al. 2022), and DICE (Sun and Li 2022). For the methods that require explicit outlier, our comparison includes OE (Hendrycks, Mazeika, and Dietterich 2019), MCD (Yu and Aizawa 2019), UDG (Yang et al. 2021), and MixOE (Zhang et al. 2023a) where all methods use remaining 800 categories of the ImageNet-1k dataset as additional outliers. Finally, we compare against various OOD synthesis methods, including VOS (Du et al. 2022), CIDER (Ming et al. 2023), NPOS (Tao et al. 2023) and DreamOOD (Du et al. 2024).
Appendix E Computational Cost
We summarize the computational cost of SONA. For outlier synthesis, generating a total of 200K images takes approximately 8.1 hours. This is significantly less time compared to other outlier-synthesis methods, as our approach does not generate outliers from random noise with full DDIM steps, but rather starts from the original image and stops early. Training a classifier with the generated outliers and an additional OE loss term takes around 5 hours. In comparison, post-hoc methods only require training a multi-class classification model on the ImageNet-200 training dataset, which takes approximately 4 hours. Our experiments were conducted using NVIDIA GeForce RTX 3090Ti GPUs.
@@ F.1 Other OOD Detection Benchmarks: CIFAR-100
CIFAR-100 contains 100 different classes, with each class containing 500 images. We use TIN (Liang and Li 2017) as the near-OOD dataset and MNIST (Deng 2012), SVHN (Netzer et al. 2011), Texture (Geirhos et al. 2019), and Place365 (Zhou et al. 2017) as far-OOD datasets. For the SONA outlier generation phase, we use CIFAR-10 as the OOD label set and generate the same number of samples as in the ID training dataset. We validated using the same implementation details as described in Appendix D.
@@ Results.
F.2 Robustness OOD Detection Benchmarks: ImageNet-200 Full-spectrum
Dataset Details.
We use ImageNet-200 as our base ID dataset, and incorporate ImageNet-C (Hendrycks and Dietterich 2019), ImageNet-R (Hendrycks et al. 2021), and ImageNet-V2 (Recht et al. 2019) as test ID datasets exhibiting covariate shift. These datasets introduce various types of distributional shifts, including corruptions, rendering styles, and natural variations, while maintaining the same semantic information with the original ImageNet-200. This setup allows us to assess our framework in distinguishing between ID samples with covariate shift and true out-of-distribution samples. We validated using the same implementation details as described in Appendix D.
Results.
F.3 Fine-grained OOD Detection Benchmark: NABirds
Dataset Details.
The NABirds dataset contains over 48,000 images of 555 North American bird species for fine-grained bird species recognition. Each image is annotated with detailed species-level labels and metadata, presenting challenges such as high inter-species similarity and intra-species variability. Following (Zhang et al. 2023a), we split the dataset into three parts. In each split, a portion of the classes is used as ID data, while the holdout classes from that split are evaluated as fine OOD data. Additionally, we utilize datasets from other sources (Maji et al. 2013; Krause et al. 2013; Chen et al. 2018) as coarse OOD data. We validated using the same implementation details as described in Appendix D.
Results.
@@
SONA (Ours) 56.26$\pm$0.26 60.17$\pm$0.13 58.22$\pm$0.19 70.42$\pm$0.33 69.24$\pm$0.24 72.43$\pm$0.18 70.69$\pm$0.25 44.52$\pm$0.16
Table 7: AUROC (%) of fine-grained OOD detection, NABirds.
@@ G.2 Data-Centric Approach
Table 9: AUROC (%) comparison with recent Data-Centric methods
@@
Appendix I Generated SONA samples
Figure 8: Class-specific visual examples of the SONA samples.
Figure 9: Class-specific visual examples of the SONA samples.

### v2

# Diffusion-based Semantic Outlier Generation via Nuisance Awareness for Out-of-Distribution Detection

@@ 1 Introduction
Figure 2: Performance comparison of Near- vs Far-OOD detection with auxiliary outliers. Only OE utilizes real outliers; the rest methods use synthetic. Outliers for the junco bird species (ID) are shown on the left.
@@ 3.4 SONA Guidance
In this section, we introduce SONA guidance, a novel approach that enables fine-grained control on each semantic and nuisance region for outlier generation. Our method draws inspiration from recent advances in image editing, which aims to completely replace target semantics. However, as they are not inherently designed for OOD detection, their application has shown limited impact in this context (see Appendix). We propose a developed approach adequate for outlier generation, allowing precise control over both $M_{\mathtt{S}}$ and $M_{\mathtt{N}}$ to prioritize semantic differences. The SONA guidance term, $\Delta_{\texttt{SONA}}$, is composed of three components as follows, each of which is described in the following paragraphs:
@@
Finally, we obtain the outliers by denoising $\mathbf{z}_{t}$ for every timestep $t=\tilde{T},\dots,1$ with our SONA guidance. To generate the pixel-space outlier image $\mathbf{x}_{\texttt{OOD}}$, we pass the denoised latent representation $\mathbf{z}_{0}$ through the decoder of the diffusion model. A detailed pseudo-code implementation of our framework can be found in the Appendix.
@@ 3.5 OOD Detection with SONA Outliers
During the test phase, we utilize the energy score (Liu et al. 2020), which effectively addresses overconfidence issues in OOD detection by assigning lower energy values to ID samples and higher energy values to OOD samples.
@@ 4.1 Experimental Setup
We mainly evaluate our framework on ImageNet-200 (Zhang et al. 2023b) as ID, a subset of 200 categories from ImageNet-1k (Deng et al. 2009). Our evaluation covers both far-OOD and challenging near-OOD scenarios. For far-OOD detection, we employ widely-used datasets such as iNaturalist, Texture, and OpenImage-O. For near-OOD detection and SSB-hard and NINCO for near-OOD detection., which have no class overlap but show close semantic similarity with ImageNet-1K.
@@
Our implementation is based on Stable Diffusion v2-base model  [^1] See https://huggingface.co/stabilityai/stable-diffusion-2-base for more details, with sampling hyperparameters consistent with (Brack et al. 2024). For selecting $\mathbf{c}_{\texttt{OOD}}$, we utilize labels from the remaining 800 classes of ImageNet-1k (disjoint classes from ImageNet-200) for the main table. The number of generated samples is equal to the number of entries in the ID training dataset. For the OOD detector training, we employ ResNet-18 (He et al. 2016) as the network architecture. Additional implementation details can be found in the Appendix.
@@ 4.2 Main Results
Table 1 presents a comprehensive comparison of OOD detection performance on ImageNet-200, highlighting the superiority of our framework. Our proposed method consistently outperforms in both near and far-OOD scenarios, affirming its effectiveness. It surpasses all benchmarks in the near-OOD setting, achieves top results in two out of three far-OOD datasets, and attains the second-best performance in the remaining one. Remarkably, our method achieves a state-of-the-art (SOTA) AUROC of 88.4% on near-OOD detection, showing exceptional performance in this challenging scenario. This score represents a notable improvement, outperforming other synthesis-based methods by 6%. Furthermore, while other baseline methods experience a sharp decline of 10% in near-OOD compared to far-OOD, our approach significantly reduces this performance gap. Another noteworthy point is we even surpass real-world outlier based methods, without requiring any additional datasets. This indicates that our methodology is also cost-effectively feasible for real-world applications. Another superior performance of SONA across a broader range of settings, including full-spectrum (e.g., covariate-shifted ID) and fine-grained benchmarks, can be found in the Appendix.
@@
Figure 6: Performance comparison of SONA guidance and global guidance across $\tilde{T}$ on ImageNet200 (ID). (a) AUROC (%) analysis. (b) LPIPS score analysis.
Figure 7: (a) Ablation study on region masking threshold $\lambda$ on ImageNet200 (ID) (b) Ablation study of SONA guidance scale $s$.
@@ 4.3 Analysis
Table 2: Comparison of SONA with different OOD prompt selection.
@@
Table 3: Ablation study on $\Delta_{\texttt{SONA}}$ components.
@@ 6 Conclusion
Acknowledgements
This work is fully supported by LG AI Research.
@@ Acknowledgements
Diffusion-based Semantic Outlier Generation via Nuisance Awareness
@@ A.2 Classifier-Free Diffusion Guidance
Appendix B Dataset Details
B.1 Near-OOD Detection
  • SSB-hard (Vaze et al. 2022) (Semantic Shift Benchmark - Hard): SSB-hard is a subset derived from the Semantic Shift Benchmark (SSB), designed to evaluate model performance under subtle semantic shifts. It contains images closely similar to ImageNet-1K, making it particularly challenging. The dataset comprises 980 classes from ImageNet-21K, sharing significant visual and semantic characteristics with ImageNet-1K.
B.2 Far-OOD Detection
  • Textures (Cimpoi et al. 2014): The Textures dataset contains images representing a wide variety of real-world patterns and surface textures, differing from the object-centric images in ImageNet-1K. It tests the model’s ability to recognize OOD samples based on textural differences, crucial for robust OOD detection.
Appendix C Implementation Details of SONA Framework
Implementation details.
Our implementation for SONA is based on the Stable Diffusion v2-base model. For more information, please refer to Footnote 1. For our best results, we set the total number of diffusion steps $T$ to 50, and at that point randomly selected a stop timestep $\tilde{T}$ from the range between 1 and 50. For the OOD text condition, $\textbf{c}_{\texttt{OOD}}$, we use text prompts based on ImageNet-800 words, which do not overlap with ImageNet-200. We set the guidance scale $s$ to 10 and the guidance threshold $\lambda$ to 0.2 for all datasets. We follow the additional hyperparameters for image sampling provided by (Brack et al. 2024). Since we use the ID training dataset directly for SONA generation, the number of SONA is equal to the number of entries in the ID training dataset. While synthesizing our SONA samples, we use 8 A40 gpus. Implementation of SONA is summarized as following algorithm 2:
Baselines.
We compare our framework against existing methods divided into three categories: Post-hoc detection methods that do not require further training, training with real-world outlier methods that require explicit outliers, and outlier synthesis-based methods. For the post-hoc baselines, we compare against OpenMax , MSP, ODIN, EBO, OpenGAN , ReAct, KNN , and DICE. For the methods that require explicit outlier, our comparison includes OE, MCD, UDG, and MixOE where all methods use remaining 800 categories of the ImageNet-1k dataset as additional outliers. Finally, we compare against various OOD synthesis methods, including VOS, CIDER , NPOS and DreamOOD.
Appendix D Computational Cost
We summarize the computational cost of SONA. For outlier synthesis, generating a total of 200K images takes approximately 8.1 hours. This is significantly less time compared to other outlier-synthesis methods, as our approach does not generate outliers from random noise with full DDIM steps, but rather starts from the original image and stops early. Additionally, SONA demonstrates superior computational efficiency compared to methods that calculate perceptual distances between input images and their diffusion-generated counterparts (Liu et al. 2023; Graham et al. 2023; Gao et al. 2023). During inference, SONA is substantially faster than these methods, which require computationally intensive diffusion reverse processes with multiple backward steps (e.g., 80x faster vs. DiffGuard (Gao et al. 2023)). This efficiency is particularly advantageous for continuously tracking and evaluating multiple datasets. Furthermore, SONA can operate effectively in on-device deployments where large diffusion models are impractical.
For classifier training, our approach requires approximately 5 hours when training from scratch with the generated outliers and an additional OE loss term. In comparison, post-hoc methods take approximately 4 hours to train a multi-class classification model on the ImageNet-200 training dataset. For ImageNet-1K, we leverage a pre-trained classifier and simply fine-tune it for SONA within a few epochs, taking only about 1 hour on 8 A100 GPUs while achieving superior performance. This efficient fine-tuning process effectively addresses concerns about classifier training costs. Our experiments were conducted using NVIDIA GeForce RTX 3090Ti GPUs.
Appendix E Full Algorithm of SONA
@@ F.1 Other OOD Detection Benchmarks: CIFAR-100
CIFAR-100 contains 100 different classes, with each class containing 500 images. We use TIN as the near-OOD dataset and MNIST, SVHN, Texture, and Place365 as far-OOD datasets. For the SONA outlier generation phase, we use CIFAR-10 as the OOD label set and generate the same number of samples as in the ID training dataset. We validated using the same implementation details as described in Appendix C.
@@ Results.
F.2 Other OOD Detection Benchmarks: ImageNet-1k
We extend the evaluation of our framework to another large-scale benchmark, the ImageNet-1k  (Krizhevsky, Hinton et al. 2009) dataset as ID, to demonstrate the further application of our framework.
Dataset Details.
ImageNet-1K contains 1,000 different classes, with approximately 1.2 million training images. It includes more high-resolution and diverse classes than ImageNet-200, making it suitable for large-scale evaluations. We use the same far-OOD and near-OOD datasets as in the ImageNet-200 experiments. For the SONA outlier generation phase, we adopted LAION dataset for OOD prompt selection that are semantically distinct from the original ImageNet classes, ensuring effective outlier generation.
Method Near-OOD Far-OOD ID ACC
SSB-hard NINCO Avg iNaturalist Textures OpenImage-O Avg
OpenMax 71.37 78.17 74.77 92.05 88.10 87.62 89.26 76.18
MSP 72.09 79.95 76.02 88.41 82.43 84.86 85.23 76.18
ODIN 71.74 77.77 74.75 91.17 89.00 88.23 89.47 76.18
EBO 72.08 79.70 75.89 90.63 88.70 89.06 89.47 76.18
VOS 76.26 80.25 78.26 92.34 91.13 90.17 91.21 77.13
OE 78.60 86.70 82.65 90.76 89.45 89.01 89.74 76.02
MixOE 73.32 82.66 77.99 91.25 89.91 88.67 89.94 75.93
DreamOOD 72.07 81.98 77.03 93.18 92.53 91.78 92.50 75.43
SONA (Ours) 82.21$\pm$0.15 88.43$\pm$0.21 85.32$\pm$0.18 95.56$\pm$0.14 93.21$\pm$0.28 95.69$\pm$0.16 94.82$\pm$0.19 76.94 $\pm$0.23
Table 6: AUROC (%) comparison of various OOD detection methods on ImageNet-1k as the ID dataset.
Results.
Table 6 presents a comprehensive comparison of OOD detection performance on ImageNet-1K. SONA demonstrates significant improvements in AUROC across both near- and far-OOD settings compared to existing baselines. The effectiveness of our framework extends to large-scale evaluations, with SONA consistently generating high-quality corrupted outliers even when using prompts from the LAION dataset. These results validate our method’s robustness across different OOD prompt sources and its scalability to comprehensive datasets like ImageNet-1K.
F.3 Robustness OOD Detection Benchmarks: ImageNet-200 Full-spectrum
Dataset Details.
We use ImageNet-200 as our base ID dataset, and incorporate ImageNet-C (Hendrycks and Dietterich 2019), ImageNet-R (Hendrycks et al. 2021), and ImageNet-V2 (Recht et al. 2019) as test ID datasets exhibiting covariate shift. These datasets introduce various types of distributional shifts, including corruptions, rendering styles, and natural variations, while maintaining the same semantic information with the original ImageNet-200. This setup allows us to assess our framework in distinguishing between ID samples with covariate shift and true out-of-distribution samples. We validated using the same implementation details as described in Appendix C.
Results.
@@
SONA (Ours) 56.26$\pm$0.26 60.17$\pm$0.13 58.22$\pm$0.20 70.42$\pm$0.33 69.24$\pm$0.24 72.43$\pm$0.18 70.69$\pm$0.25 44.52$\pm$0.16
F.4 Fine-grained OOD Detection Benchmark: NABirds
Dataset Details.
The NABirds dataset contains over 48,000 images of 555 North American bird species for fine-grained bird species recognition. Each image is annotated with detailed species-level labels and metadata, presenting challenges such as high inter-species similarity and intra-species variability. Following (Zhang et al. 2023a), we split the dataset into three parts. In each split, a portion of the classes is used as ID data, while the holdout classes from that split are evaluated as fine OOD data. Additionally, we utilize datasets from other sources (Maji et al. 2013; Krause et al. 2013; Chen et al. 2018) as coarse OOD data. We validated using the same implementation details as described in Appendix C.
Results.
@@
Table 8: AUROC (%) of fine-grained OOD detection, NABirds.
@@ G.2 Data-Centric Approach
Table 10: AUROC (%) comparison with recent Data-Centric methods
@@
Appendix I Challenging Cases
We present some challenging cases, including geological formations (e.g., cliffs, lakeside, iceberg, sand dune), to demonstrate the capabilities of SONA samples. These geological formations exhibit diverse and complex textures, characterized by intricate and detailed patterns, making structural delineation challenging. For instance, cliffs feature irregular, rugged surfaces with layered compositions, while lakesides involve dynamic interactions between water and surrounding vegetation. Similarly, brain corals display highly intricate, maze-like patterns that require precise semantic identification. Our primary goal is not to achieve perfect semantic replacement but to identify and guide the semantics to generate natural corruptions. While target semantic masks and stop timesteps varied across cases, SONA consistently produced realistic outliers while maintaining contextual coherence. This highlights SONA’s ability to robustly identify even challenging semantics, demonstrating its adaptability across diverse and intricate scenarios.
Figure 8: Challenging case of the SONA samples on ImageNet-1k dataset as the ID dataset.
Appendix J More Qualitative Results
Figure 9: Class-specific visual examples of the SONA samples on ImageNet-200 dataset as the ID dataset.
Figure 10: Class-specific visual examples of the SONA samples on ImageNet-200 dataset as the ID dataset.